A meeting of consuls was attended by 4 Chinese and 2 Japanese. If three consuls where selected at random one after the other, determine the values of the random variable J representing the number of Japanese. 


1
Expert's answer
2022-02-24T06:30:17-0500
"4+2=6"

There are "\dbinom{6}{3}=20" possible outcomes.

The possible values of "J" are "0,1,2."


"P(J=0)=\dfrac{\dbinom{2}{0}\dbinom{4}{3-0}}{\dbinom{6}{3}}=\dfrac{1(4)}{20}=\dfrac{1}{5}"

"P(J=1)=\dfrac{\dbinom{2}{1}\dbinom{4}{3-1}}{\dbinom{6}{3}}=\dfrac{2(6)}{20}=\dfrac{3}{5}"


"P(J=2)=\dfrac{\dbinom{2}{2}\dbinom{4}{3-2}}{\dbinom{10}{4}}=\dfrac{1(4)}{20}=\dfrac{1}{5}"


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 j & 0 & 1 & 2 \\ \hline
 p(j) & 1/5 & 3/5 & 1/5 
\end{array}"